OpenStudy (anonymous):

Need help. I can not figure out how to factor this expression to find the vertex. Also if you could show me how you find the answer that would be awesome. @Nurali

OpenStudy (anonymous):

f(x) = -16x2 + 8x + 3

OpenStudy (ranga):

f(x) = -16x2 + 8x + 3 = -16(x^2 - x/2 -3/16) = -16{ (x-1/4)^2 - 1/16 - 3/16 } simplify.

OpenStudy (anonymous):

is that divided by? sorry math not my strong suit.

OpenStudy (ranga):

f(x) = -16x^2 + 8x + 3 Factor -16 out. f(x) = -16(x^2 - x/2 -3/16) complete the square: f(x) = -16{ (x-1/4)^2 - (1/4)^2 - 3/16 } = -16{ (x-1/4)^2 - 1/16 - 3/16 } = -16{ x-1/4)^2 - 4/16} = -16{ x-1/4)^2 - 1/4} f(x) = -16(x - 1/4)^2 + 4 Vertex at (1/4, 4)

OpenStudy (anonymous):

Okay. Thank you.